# Differentiable Structure Learning for General Binary Data

### v1

@@ 1 Introduction
Outside the differentiable paradigm, decades of work have produced both score-based and constraint-based algorithms for discrete causal discovery. Score-based searches (e.g., greedy equivalence search (Chickering, 2002)) perform well on small graphs, but their greedy, local exploration of a super-exponential DAG space often traps them in local optima and limits scalability. Constraint-based procedures—such as the PC algorithm and its variants (Spirtes et al., 2000; Cui et al., 2016)—rely on conditional-independence tests that become unreliable with finite samples or measurement noise. Many discrete methods also impose additional simplifications (additive-noise models (Peters et al., 2010), hidden compact representations (Cai et al., 2018), linear effects (Li and Shimizu, 2018), latent Gaussian variables (Andrews et al., 2019)), which can lead to model misspecification. Even the most recent differentiable approaches evaluated on discrete data lack a formal justification. For instance, Bello et al. (2022) apply continuous optimization methods to discrete data without fully accounting for discrete-specific characteristics, introducing implicit assumptions and evaluation metrics. These observations underscore the need for a general, theoretically-sound framework for differentiable structure learning in discrete data.
@@
    We start with general binary data, without assuming a particular data generating process, and use the multivariate Bernoulli distribution (Definition 1) as a general representation for any such model. We prove that, in this fully general setting, the underlying DAG is non-identifiable and show all the graph–parameter pairs compatible with the observed distribution (Theorem 1).
@@
    We formulate learning of the sparsest graph and its parameters as a single differentiable optimization (13). We establish that any global minimizer yields the sparsest valid DAG, and that under the faithfulness assumption all such DAGs are within in the same Markov equivalence class (Theorem 3). Moreover, based on our framework, we provide theoretical justification for prior works (Deng et al., 2023a; Bello et al., 2022).
@@
    We conduct experiments demonstrating that our method can capture the most general causal relationships. Furthermore, we introduce a two‐stage approach, Notears-MLP-Reg, which reduces computational complexity and outperforms existing baselines.
@@ 2 Related work
Classical causal‐discovery methods for discrete data fall into two main categories. Constraint‐based algorithms (e.g. PC (Spirtes and Glymour, 1991), MMHC (Tsamardinos et al., 2003), Copula‐PC (Cui et al., 2016)) use conditional‐independence tests to infer edges, but finite samples and measurement noise often render these tests unreliable. Score‐based approaches assign each candidate DAG a goodness‐of‐fit score (e.g. BIC, BDeu (Maxwell Chickering and Heckerman, 1997; Heckerman et al., 1995), generalized scores (Huang et al., 2018; Li and Shimizu, 2018)) and then search for the optimal structure. Greedy strategies such as GES (Chickering, 2002) can work on small graphs but suffer from NP hardness (Chickering, 1996; Chickering et al., 2004) and can become trapped in local optima (Lam et al., 2022; Andrews et al., 2023). Both paradigms typically rely on strong parametric assumptions—additive‐noise models (Peters et al., 2010), hidden compact representations (Cai et al., 2018), linear effects (Li and Shimizu, 2018), latent Gaussian variables (Andrews et al., 2019)—which may not hold in practice. A few methods target bivariate causal discovery (Wei et al., 2022; Liu and Chan, 2016; Budhathoki and Vreeken, 2018), but they do not fit to multivariate problem common in modern applications. In summary, existing discrete causal‐discovery techniques either lack statistical robustness in finite samples or impose restrictive assumptions, motivating the need for scalable, assumption‐light methods that handle high‐dimensional binary data.
@@
Since the introduction of a smooth characterization of acyclicity by Zheng et al. (2018), which allows treating the adjacency matrix as a continuous optimization variable, a large body of work has emerged in differentiable structure learning. Subsequent research has extended this framework in several directions: toward nonlinear models (Zheng et al., 2020; Yu et al., 2019; Lachapelle et al., 2019; Zhu et al., 2019), theoretical analysis from an optimization perspective (Wei et al., 2020; Ng et al., 2022; Deng et al., 2023b), designing more stable and computationally efficient acyclicity constraints (Bello et al., 2022; Nazaret et al., 2023; Yu et al., 2019; Zhang et al., 2022, 2025), understanding score function properties (Ng et al., 2020; Deng et al., 2024; Seng et al., 2024), and incorporating extra side information (Ban et al., 2024). Despite this progress, most existing methods are fundamentally designed for continuous variables and do not address the specific challenges of discrete data. The common reliance on Gaussianity or continuous-variable assumptions—leading to objectives like mean-squared error or Gaussian log-likelihood—introduces potential biases and inconsistencies when applied to discrete data. Thus, naively transferring continuous-data DAG learning methods to discrete domains is problematic and necessitates principled modifications.
@@
Only a handful of recent studies have extended differentiable structure learning to discrete data by imposing specific parametric forms on the conditionals. For example, most works (Zeng et al., 2022; Bello et al., 2022; Deng et al., 2023a) assume a linear SEM in which each discrete node’s probability is given by a logistic link on a linear combination of its parents. Such restrictive assumptions can fail to capture the rich, higher‐order dependencies present in general discrete distributions. Another line of work in differentiable structure learning studies nonlinear models via neural networks (Zheng et al., 2020; Lachapelle et al., 2019; Zhu et al., 2019), which in principle could accommodate discrete inputs. However, these methods treat inputs as generic real‐valued vectors and do not formally address the unique features of discrete data, neither in theory nor in empirical evaluation. A recent contribution by Deng et al. (2024) considers broad parametric families that include general binary distributions, but it remains unclear how to extend their approach to the discrete variables. Collectively, these limitations motivate the development of a truly general differentiable framework capable of learning arbitrary discrete causal relationships without relying on narrow parametric assumptions.
@@ 3 Preliminaries
Let $G=(V,E)$ denote a directed graph on $p$ nodes, with vertex set $V=[p]\coloneq\{1,\ldots,p\}$ and edge set $E\subset V\times V$, where $(i,j)\in E$ indicates the presence of a directed edge from node $i$ to node $j$. We associate each node $i\in V$ to a random variable $X_{i}$, and let $X=(X_{1},\ldots,X_{p})$. In this work we focus on discrete random variables. Any variable with $T$ categories can be represented by (T1)(T-1) binary indicators (Wei et al., 2018). For instance, if $Z\in\{0,1,2\}$, then introduce $Z^{k}\coloneqq\mathbbm{1}\{Z=k\},\;k=0,1,2$, indicating whether $Z=k$ for $k=0,1,2$. so that $Z$ is replaced by three binary indicators. Hence, without loss of generality, we only consider binary data in the sequel. Extending the results to general discrete data only need additional notation and does not require new technical ideas.
@@
$\begin{split}P(X=x)=\prod_{S\subseteq[p]}p(1_{S})^{\,\prod_{j\in S}x_{j}\;\prod_{j\notin S}(1-x_{j})}\end{split}$ (1)
where $1_{S}\in\{0,1\}^{p}$ for the vector with ones on $S$ and zeros elsewhere and each entry of $\bm{p}=\bigl(p(1_{S})\bigr)_{S\subseteq[p]}\in\mathbb{R}^{2^{p}}$ is the probability mass of a distinct configuration. Such $\bm{p}$ is called the general parameter. Let $\mathbf{1}_{2^{p}}\in\mathbb{R}^{2^{p}}$ denote the all-ones vector; normalization requires $\mathbf{1}_{2^{p}}^{\top}\bm{p}=1$.
It is hard to figure out the dependence between each component based on the form of probability mass function in (1). As a consequence, define $B^{S}(x)=\prod_{j\in S}x_{j}$ for any $S\subseteq[p]$ with convention $B^{\emptyset}=1$. Rewrite (1) in exponential–family form:
@@
where the coefficient $\{f_{S}\}$ are the natural parameters. The corresponding natural‐parameter vector is $\bm{f}=(f^{0},f^{1},\ldots,f^{p},\ldots,f^{1\ldots p})\in\mathbb{R}^{2^{p}}$. Throughout, the superscript set $S$ in $f^{S}$ is treated as unordered; e.g. for $S=\{1,2\}$ we have $f^{12}=f^{21}$.
Because the MVB is an exponential family, there is a one-to-one correspondence between the natural parameters $\bm{f}$ and the general parameter $\bm{p}$; explicit conversion formulas are provided in Appendix C.1. Expressing the model as in (2) will later allow us to identify variable dependencies directly from the coefficients $\bm{f}$ when reconstructing causal graphs. A simple but useful closure property follows immediately.
@@ 4 General discrete data: a nonidentifiable model
We first establish that, for fully general binary data, neither the causal structure nor the associated parameters can be identified from a single observational environment  [^1] By “single environment” we mean i.i.d. samples from a fixed observational distribution $p(X)$ with no interventions or exogenous shifts. In contrast, with multiple observational environments—or with interventional data—causal structure can be identified under standard invariance assumptions; see Guo et al. (2022); Huang et al. (2020).. The intuition is simple: for any fixed topological sort (or order) of the variables (Definition 2) there exists a unique causal structure and associated parameters that reproduces the observed distribution exactly. Because a set of $p$ variables admits up to $p!$ distinct topological orders, it renders the model fundamentally non-identifiable.
4.1 Conditional distribution: higher order interaction
Definition 2 (Topological sort).
Let $G=(V,E)$ be a directed graph with vertex set $V=[p]$. A topological sort is a permutation $\pi$ of $V$ such that $X_{\pi(i)}\to X_{\pi(j)}\;\Longrightarrow\;i<j.$ Equivalently, the permutation orders the nodes so that every directed edge points from an earlier to a later position. A directed graph is acyclic iff at least one such ordering exists; the ordering need not be unique.
Given a permutation $\pi$, the joint distribution admits the standard factorization
$P(X)=\prod_{j=1}^{p}P(X_{\pi(j)}\mid X_{\pi(1),\ldots,\pi(j-1)})$
Let us use the joint law $P(X)$ given by the exponential form in (2). We examine the conditional distribution $P\!\bigl(X_{\pi(j)}\,\big|\,X_{\pi(1)},\dots,X_{\pi(j-1)}\bigr)$. Without loss of generality let $\pi=(1,\dots,p)$ and focus on $j=p$; other choices of $\pi$ and $j$ lead to the same algebra with heavier notation. Then
$\begin{split}&P(X_{p}=1\mid X_{1}=x_{1},\ldots,X_{p-1}=x_{p-1})\\ =&\sigma\left(\sum_{r=1}^{p}\left(\sum_{1\leq j_{1}<j_{2}<\ldots<j_{r}=p\leq p}f^{j_{1}\ldots j_{r-1}p}x_{j_{1}}\ldots x_{j_{r-1}}x_{p}\right)\right)\\ =&\sigma\left(f^{p}+f^{1p}x_{1}+f^{2p}x_{2}\ldots f^{p-1,p}f_{p-1}+f^{12p}x_{1}x_{2}+\ldots+f^{1\ldots p}x_{1}\ldots x_{p-1}\right)\end{split}$ (3)
where $\sigma(z)=1/\left(1+\exp(-z)\right)$. The full derivation appears in Appendix C.2. Equation (3) highlights that, for general binary data, all higher order interaction terms are present (Dai et al., 2013; Enouen and Sugiyama, 2024). Omitting these higher-order interactions—e.g., restricting attention to first-order (additive) effects (Ye et al., 2024; Zeng et al., 2022)—can misrepresent real data.
@@ 4.2 Nonidentifiability and Equivalence
[…]
@@
$\Phi(X)=\bigl[1,\;X_{1},\dots,X_{p},\;X_{1}X_{2},X_{1}X_{3},\dots,X_{p-1}X_{p},\;X_{1}X_{2}X_{3},\dots,X_{1}\cdots X_{p}\bigr].$ (5)
For example, when $p=3$ and $X=(X_{1},X_{2},X_{3})$, then the extended feature vector
$\Phi(X)=[1,X_{1},X_{2},X_{3},X_{1}X_{2},X_{1}X_{3},X_{2}X_{3},X_{1}X_{2}X_{3}]\in\mathbb{R}^{8}$
When $\Phi$ is applied to data matrix ${\mathbf{X}}\in\mathbb{R}^{n\times p}$, it operates row-wise.
@@
$\bm{f}_{\pi,j}=[\underbrace{f_{\pi,j}^{\pi(j)}}_{\text{constant}},\underbrace{f_{\pi,j}^{\pi(1)\pi(j)}}_{X_{\pi(1)}},\underbrace{f_{\pi,j}^{\pi(2)\pi(j)}}_{X_{\pi(2)}},\ldots,\underbrace{f_{\pi,j}^{\pi(j-1)\pi(j)}}_{X_{\pi(j-1)}},\underbrace{f_{\pi,j}^{\pi(1)\pi(2)\pi(j)}}_{X_{\pi(1)}X_{\pi(2)}},\ldots,\underbrace{f_{\pi,j}^{\pi(1)\ldots\pi(j-1)\pi(j)}}_{X_{\pi(1)}\ldots X_{\pi(j-1)}}].$ (6)
@@
With these definitions in place we can now recover the full causal structure and its parameters. Fix an ordering $\pi$ and consider any $j\in[p]$. By Corollary 1, both the marginal $P(X_{\pi(j)},X_{\pi(1)},\ldots,X_{\pi(j-1)})$ and conditional $P(X_{\pi(j)}\mid X_{\pi(1),\ldots,\pi(j-1)})$ remain multivariate Bernoulli. Let ${\mathbf{X}}=({\mathbf{X}}_{1},\dots,{\mathbf{X}}_{p})\in\mathbb{R}^{n\times p}$ be $n$ i.i.d. samples from $X\sim\mathrm{MultiBernoulli}(\bm{p})$, one sample per row. Hence we can apply the same logic as in Section 4.1
@@
Equation (7) is written as a form of structural equation model (SEM) [^2] Formal introduction can be founded in Appendix C.5 (Peters et al., 2017), a framework widely used for structure learning because it provides a generative description of how each variable arises from its direct causes. A central appeal of SEMs is their universality: every probability distribution can, in principle, be represented by a suitably chosen SEM (Proposition 7.1 in Peters et al., 2017). In practice, however, SEM-based methods almost always impose strong parametric forms—linear, additive, or low‐order interactions (Ye et al., 2024; Zeng et al., 2022; Gu et al., 2019; Zhao et al., 2024a)—which risk ignoring important structure in real data (Dai et al., 2013; Enouen and Sugiyama, 2024). The problem is particularly acute for discrete data, where higher-order dependencies are easily overlooked, sharply limiting the model’s expressive power. By contrast, Theorem 1 shows that general binary distributions necessarily involve higher-order interaction terms. Omitting these terms therefore precludes an exact representation and may lead to incorrect causal conclusions.
Thus, Theorem 1 indicates that, for every topological order $\pi$ we can construct a distinct SEM that reproduces the same distribution. The model is therefore non-identifiable: multiple parameter–graph pairs $(\bm{f}_{\pi},G_{\pi})$ obtained by procedure before give rise to the identical law $\mathrm{MultiBernoulli}(\bm{p})$. Collect all such pairs into the equivalence class
@@
It is well-known that causal DAGs are non-identifiable from observational data alone (Heckerman et al., 1995). However, Theorem 1 provides a novel and comprehensive characterization of all DAG–parameter pairs that exactly reproduce a given distribution within the general multivariate Bernoulli framework.

### v2

# Differentiable Structure Learning and Causal Discovery for General Binary Data

@@ 1 Introduction
Outside the differentiable paradigm, decades of work have produced both score-based (e.g. Chickering, 2002) and constraint-based (e.g. Spirtes and Glymour, 1991) algorithms for causal discovery. Many discrete methods also impose additional simplifications (additive-noise models (Peters et al., 2010), hidden compact representations (Cai et al., 2018), linear effects (Li and Shimizu, 2018), latent Gaussian variables (Andrews et al., 2019)), which may not apply to discrete data. Even the most recent differentiable approaches evaluated on discrete data lack a formal justification: For instance, Bello et al. (2022) apply continuous optimization methods to discrete data under a logistic regression model, although the underlying identifiability and theory for such a model has yet to be worked out. These observations underscore the need for a general, theoretically-sound framework for differentiable structure learning in discrete data.
@@
    We start with general binary data, without assuming a particular data generating process, and use the multivariate Bernoulli distribution (Definition 1) as a general representation for any such model. We prove that, in this fully general setting, the underlying DAG is non-identifiable and characterize which graph–parameter pairs are compatible with the observed distribution (Theorem 1).
@@
    We formulate learning of the sparsest graph and its parameters as a single differentiable program (13). We show that even without assuming faithfulness, any global minimizer yields the sparsest valid DAG (Theorem 3). Moreover, based on our framework, we provide theoretical justification for prior works (Deng et al., 2023a; Bello et al., 2022).
@@
    We conduct experiments demonstrating that our method can capture general causal relationships between discrete variables. Furthermore, we introduce a two‐stage approach, BiNOTEARS, which reduces computational complexity and outperforms existing baselines.
@@ 2 Related work
Classical causal‐discovery methods for discrete data fall into two main categories. Constraint‐based algorithms (e.g. PC (Spirtes and Glymour, 1991), MMHC (Tsamardinos et al., 2003), Copula‐PC (Cui et al., 2016)) use conditional‐independence tests to infer edges. Score‐based approaches assign each candidate DAG a goodness‐of‐fit score (e.g. BIC, BDeu (Maxwell Chickering and Heckerman, 1997; Heckerman et al., 1995), generalized scores (Huang et al., 2018; Li and Shimizu, 2018)) and then search for the optimal structure. Greedy strategies such as GES (Chickering, 2002) are known to be NP-hard (Chickering, 1996; Chickering et al., 2004) and can get trapped in local optima (Lam et al., 2022; Andrews et al., 2023). Moreover, score-based search traditionally relies on strong parametric assumptions—e.g. additive‐noise models (Peters et al., 2010), hidden compact representations (Cai et al., 2018), linear effects (Li and Shimizu, 2018), latent Gaussian variables (Andrews et al., 2019)—which may not hold in practice. A recent exception for general nonparametric models is Aragam (2024). A few methods target bivariate causal discovery (Wei et al., 2022; Liu and Chan, 2016; Budhathoki and Vreeken, 2018), but they do not fit to multivariate problems. As a result, existing discrete causal‐discovery techniques either lack statistical robustness in finite samples or impose restrictive assumptions, motivating the need for scalable, assumption‐light methods that handle high‐dimensional binary data.
@@
Since the introduction of a smooth characterization of acyclicity by Zheng et al. (2018), which allows treating the adjacency matrix as a continuous optimization variable, a large body of work has emerged in differentiable structure learning. Subsequent research has extended this framework in several directions: Nonlinear models (Zheng et al., 2020; Yu et al., 2019; Lachapelle et al., 2019; Zhu et al., 2019), optimization theory (Wei et al., 2020; Ng et al., 2022; Deng et al., 2023b), designing more stable and computationally efficient acyclicity constraints (Bello et al., 2022; Nazaret et al., 2023; Yu et al., 2019; Zhang et al., 2022, 2025), understanding score function properties (Ng et al., 2020; Deng et al., 2024; Seng et al., 2024), and incorporating extra side information (Ban et al., 2024). Despite this progress, most existing methods are fundamentally designed for continuous variables and do not address the specific challenges of discrete data. The common reliance on Gaussianity or continuous-variable assumptions—leading to objectives like mean-squared error or Gaussian log-likelihood—introduces potential biases and inconsistencies when applied to discrete data.
@@
Only a handful of recent studies have extended differentiable structure learning to discrete data by imposing specific parametric forms on the conditionals. For example, most works (Zeng et al., 2022; Bello et al., 2022; Deng et al., 2023a) assume a linear SEM in which each discrete node’s probability is given by a logistic link on a linear combination of its parents. Such restrictive assumptions can fail to capture the rich, higher‐order dependencies present in general discrete distributions. Another line of work in differentiable structure learning studies nonlinear models via neural networks (Zheng et al., 2020; Lachapelle et al., 2019; Zhu et al., 2019), which in principle could accommodate discrete inputs. However, these methods treat inputs as generic real‐valued vectors and do not formally address the unique features of discrete data, neither in theory nor in empirical evaluation. A recent contribution by Deng et al. (2024) proposes a general differentiable framework for parametric families, but the necessary details to connect generic parametrizations of discrete data to structure learning (in particular, the adjacency matrix), are missing. By contrast, one the of the key contributions in this work is to develop a tractable parametrization via the multivariate Bernoulli distribution that allows explicit representation of the adjacency matrix in a fully differentiable framework without restricting the class of discrete distributions.
@@ 3 Preliminaries
Let $G=(V,E)$ denote a directed graph on $p$ nodes, with vertex set $V=[p]\coloneq\{1,\ldots,p\}$ and edge set $E\subset V\times V$, where $(i,j)\in E$ indicates the presence of a directed edge from node $i$ to node $j$. We associate each node $i\in V$ to a random variable $X_{i}$, and let $X=(X_{1},\ldots,X_{p})$. In this work we focus on discrete random variables. Any variable with $T$ categories can be represented by (T1)(T-1) binary indicators (Wei et al., 2018). For instance, if $Z\in\{0,1,2\}$, then introduce $Z^{k}\coloneqq\mathbbm{1}\{Z=k\},\;k=0,1,2$, indicating whether $Z=k$ for $k=0,1,2$. so that $Z$ is replaced by three binary indicators. Hence, without loss of generality, we only consider binary data in the sequel. Extending the results to general discrete data only needs additional notation and does not require new technical ideas.
@@
$\begin{split}P(X=x)=\prod_{S\subseteq[p]}p(1_{S})^{\,\prod_{j\in S}x_{j}\;\prod_{j\notin S}(1-x_{j})}.\end{split}$ (1)
where $1_{S}\in\{0,1\}^{p}$ for the vector with ones on $S$ and zeros elsewhere and each entry of $\bm{p}\in\mathbb{R}^{2^{p}}$ is the probability mass of a distinct configuration. Such $\bm{p}$ is called the general parameter. Let $\mathbf{1}_{2^{p}}\in\mathbb{R}^{2^{p}}$ denote the all-ones vector; normalization requires $\mathbf{1}_{2^{p}}^{\top}\bm{p}=1$.
Based on (1), it is not clear how to determine the dependence between each variable. For this, define $B^{S}(x)=\prod_{j\in S}x_{j}$ for any $S\subseteq[p]$ with convention $B^{\emptyset}(x)=1$. Rewrite (1) in exponential–family form:
@@
where the coefficient $\{f^{S}\}$ are the natural parameters. The corresponding natural‐parameter vector is $\bm{f}=(f^{0},f^{1},\ldots,f^{p},\ldots,f^{1\ldots p})\in\mathbb{R}^{2^{p}}$. Throughout, the superscript set $S$ in $f^{S}$ is treated as unordered; e.g. for $S=\{1,2\}$ we have $f^{12}=f^{21}$. Because the MVB is an exponential family, there is a one-to-one correspondence between the natural parameters $\bm{f}$ and the general parameter $\bm{p}$; explicit conversion formulas are provided in Appendix C.1. Expressing the model as in (2) will later allow us to identify variable dependencies directly from the coefficients $\bm{f}$ when reconstructing causal graphs. A simple but useful closure property follows immediately.
@@ 4 General discrete data: A nonidentifiable model
We first establish that, for fully general binary data, neither the causal structure nor the associated parameters can be identified from a single observational environment (Guo et al., 2022; Huang et al., 2020). The intuition is simple: For any fixed topological sort [^1] A topological sort is a permutation $\pi$ of $V$ such that $X_{\pi(i)}\to X_{\pi(j)}\;\Longrightarrow\;i<j.$ of the variables there exists a unique causal structure and associated parameters that reproduces the observed distribution exactly. Since there are $p!$ distinct topological sorts, this renders the model fundamentally non-identifiable.
4.1 Conditional distributions and higher order interaction
First, we provide a useful expression for conditional distributions in this model. Given any permutation $\pi$ on $V$, the joint distribution admits the standard factorization
$P(X)=\prod_{j=1}^{p}P(X_{\pi(j)}\mid X_{\pi(1),\ldots,\pi(j-1)}).$
Let us use the joint law $P(X)$ given by the exponential form in (2). Without loss of generality let $\pi=(1,\dots,p)$ and focus on $j=p$; other choices of $\pi$ and $j$ lead to the same algebra with heavier notation. Then
$\begin{split}&\Pr(X_{p}=1\mid X_{-p}=x_{-p})\\ =&\sigma\left(f^{p}+f^{1,p}x_{1}+\ldots f^{p-1,p}x_{p-1}+f^{12,p}x_{1}x_{2}+\ldots+f^{1\ldots,p}x_{1}\ldots x_{p-1}\right)\\ =&\sigma\!\left(\sum_{S\subseteq[p-1]}f^{S,p}\;B^{S}(x)\right)\end{split}$ (3)
where $p-1]=\{1,\ldots,p-1\}$ and $\sigma(z)=1/\left(1+\exp(-z)\right)$. The full derivation appears in Appendix C.2. Equation (3) highlights that, for general binary data, all higher order interaction terms are present (Dai et al., 2013; Enouen and Sugiyama, 2024). Omitting these higher-order interactions—e.g., restricting attention to first-order (additive) effects (Ye et al., 2024; Zeng et al., 2022)—can misrepresent real data.
@@ 4.2 Nonidentifiability and Equivalence
$X_{j}\rightarrow X_{p}\Leftrightarrow\exists\;S\subseteq[p-1]\text{ with }j\in S,\text{ such that }f^{S,p}\neq 0\Leftrightarrow\sum_{S\subseteq[p-1],j\in S}\left(f^{S,p}\right)^{2}>0.$ (4)
@@
$\Phi(X)=\bigl[1,\;X_{1},\dots,X_{p},\;X_{1}X_{2},\dots,X_{p-1}X_{p},\;X_{1}X_{2}X_{3},\dots,X_{1}\cdots X_{p}\bigr].$ (5)
For example, when $p=3$ and $X=(X_{1},X_{2},X_{3})$, then the extended feature vector $\Phi(X)=[1,X_{1},X_{2},X_{3},X_{1}X_{2},X_{1}X_{3},X_{2}X_{3},X_{1}X_{2}X_{3}]\in\mathbb{R}^{8}$. When $\Phi$ is applied to data matrix ${\mathbf{X}}\in\mathbb{R}^{n\times p}$, it operates row-wise.
@@
$\bm{f}_{\pi,j}=[\underbrace{f_{\pi,j}^{\pi(j)}}_{\text{constant}},\underbrace{f_{\pi,j}^{\pi(1),\pi(j)}}_{X_{\pi(1)}},\underbrace{f_{\pi,j}^{\pi(2),\pi(j)}}_{X_{\pi(2)}},\ldots,\underbrace{f_{\pi,j}^{\pi(j-1),\pi(j)}}_{X_{\pi(j-1)}},\underbrace{f_{\pi,j}^{\pi(1)\pi(2),\pi(j)}}_{X_{\pi(1)}X_{\pi(2)}},\ldots,\underbrace{f_{\pi,j}^{\pi(1)\ldots\pi(j-1),\pi(j)}}_{X_{\pi(1)}\ldots X_{\pi(j-1)}}].$ (6)
@@
With these definitions in place we can now recover the full causal structure and its parameters. Fix an ordering $\pi$ and consider any $j\in[p]$. By Corollary 1, both the marginal $P(X_{\pi(j)},X_{\pi(1)},\ldots,X_{\pi(j-1)})$ and conditional $P(X_{\pi(j)}\mid X_{\pi(1),\ldots,\pi(j-1)})$ remain multivariate Bernoulli. Let ${\mathbf{X}}=({\mathbf{X}}_{1},\dots,{\mathbf{X}}_{p})\in\mathbb{R}^{n\times p}$ be $n$ i.i.d. samples from $X\sim\mathrm{MultiBernoulli}(\bm{p})$, one sample per row. Hence we can apply the same logic as in Section 4.1:
@@
Equation (7) expresses a general discrete model as a structural equation model (SEM) [^2] Formal details can be found in Appendix C.5, a framework widely used for structure learning because it provides a generative description of how each variable arises from its direct causes. A central appeal of SEMs is their universality: Every probability distribution can, in principle, be represented by a suitably chosen SEM (Proposition 7.1 in Peters et al., 2017). In practice, however, SEM-based methods almost always impose strong parametric forms—linear, additive, or low‐order interactions (Ye et al., 2024; Zeng et al., 2022; Gu et al., 2019; Zhao et al., 2024a)—which risk ignoring important structure in real data (Dai et al., 2013; Enouen and Sugiyama, 2024). The problem is particularly acute for discrete data, where higher-order dependencies are easily overlooked, sharply limiting the model’s expressive power. By contrast, Theorem 1 shows that general binary distributions necessarily involve higher-order interaction terms. Omitting these terms therefore precludes an exact representation and may lead to incorrect causal conclusions.
Thus, the Theorem 1 indicates that, for every topological order $\pi$ we can construct a distinct SEM that reproduces the same distribution. The model is therefore non-identifiable: multiple parameter–graph pairs $(\bm{f}_{\pi},G_{\pi})$ obtained by procedure before give rise to the identical law $\mathrm{MultiBernoulli}(\bm{p})$. Collect all such pairs into the equivalence class
@@
It is well-known that causal DAGs are non-identifiable from observational data alone (Heckerman et al., 1995; Guo et al., 2022; Huang et al., 2020). However, Theorem 1 provides a novel and comprehensive characterization of all DAG-parameter pairs that exactly reproduce a given distribution within the general multivariate Bernoulli framework.